OpenStudy (the_legend_of_kris):

Sean is taking some friends on a picnic. They'll need to follow a path to get to the picnic spot. A map of the path is based on a scale of 1:30,000, in cm. - If the path is 20cm on the map, how many kilometers is the actual path

OpenStudy (lapinblanc):

Ratio solutions follow: a/b=c/d you have 1/30,000=20/x solve for x Convert to km by multiplying by 1x10^-5

OpenStudy (lapinblanc):

1/30000=20/x multiply by reciprocal 30,000(1/30,000)=(20/x)30,000 1=600,000/x x*1=(600,000/x)x x=600,000cm Km conversion to cm is 1km=1x10^-5 1x10^-5 is the same as 0.00001 So 600,000 x 0.0001=6km

OpenStudy (lapinblanc):

I'm sorry, to be clear, conversion was 1cm=1x10^5km. Everything else is accurate.